OpenStudy (anonymous):

Tracy ran for seven-twelfths hours and Reina ran for one-third hours. How many more hours did Tracy run

OpenStudy (anonymous):

Tracy ran for 7/12 hours. Reina ran for 1/3 hours. It's good to make them have equal denominators: 1/3 = (1*4)/(3*4)= 4/12 Now, subtract Tracy's from Reina's: 7/12 - 4/12 = 3/12 = 1/4 So, Tracy ran 1/4 hours more than Reina.